# Emotion-LLaMAv2 and MMEVerse: A New Framework and Benchmark for Multimodal Emotion Understanding

### v1

@@ 5.3 Main Results
TABLE VI: The Emotion-LLaMAv2 performance trained on different MER instruction tuning datasets.
@@
TABLE VII: Ablation study of the Conv-Attention module.
@@
TABLE VIII: Comparison of different pre-fusion architectures.
@@
TABLE IX: Comparison of different training schemes.
@@ 5.4 Data Evaluation
Effectiveness of MMEVerse. Table VI presents an evaluation of recent multimodal emotion instruction tuning datasets, including MERR-Coarse, MERR-Fine, and MER-Caption+. We train Emotion-LLaMAv2 on these datasets with their emotion categories and showcase its performance on both MER-UniBench and MMEVerse-Bench. The results reveal that training on MMEVerse-Train yields significantly better performance, particularly on MMEVerse-Bench. This can be attributed to two main factors: its data diversity and annotation reliability. On the one hand, its diverse sources, spanning from YouTube videos to interviews, capture more authentic "in-the-wild" emotional expressions, equipping the model with superior generalization capabilities. On the other hand, MMEVerse is more accurate in annotation quality since each emotion label is manually annotated, enabling the model to learn true multimodal-emotion correlations without label noise.
@@ 5.5 Ablation Studies
Effectiveness of the Conv-Attention Module. Table VII reports the evaluation of the Conv-Attention module. It significantly improves Stage 1 performance, increasing the average score by 1.48% on MER-UniBench and 0.92% on MMEVerse-Bench. This highlights its importance: pre-fusion representations complement unimodal embeddings through interaction before feeding into the LLM. For models trained in Stage 2, the gains remain but are slightly reduced, as complex reasoning training can offset some improvements.
To further assess the pre-fusion module, we compare our Conv-Attn architecture against the Q-Former and the MLP-based attention mechanism from AffectGPT [40]. As shown in Table VIII, the inclusion of any pre-fusion module consistently improves performance over the baseline, validating the strategy of refining modal features prior to joint processing by the LLM. Among individual architectures, both the Q-Former and our Conv model perform strongly, suggesting that preserving local temporal information during pre-fusion is particularly effective. Ultimately, Conv-Attn achieves the highest performance, reflecting a synergistic effect between its components: convolutional layers excel at extracting local temporal features, while attention mechanisms capture long-range dependencies and global context. By integrating these complementary strengths, the model produces richer and more comprehensive representations, leading to superior final performance.
Effectiveness of the perception-to-cognition training scheme. Table IX presents an evaluation of different strategies for instruction tuning. While one might argue that a joint training scheme for recognition and reasoning is simpler and more efficient, akin to our preliminary Emotion-LLaMA, surprisingly, this straightforward approach yields significantly poorer performance on both benchmarks listed in Table IX. This can be attributed to the difficulty of learning both perception and cognition abilities from scratch, in contrast to a curriculum learning approach.
TABLE X: Ablation study on Modality Encoders. All the models are trained on MMEVerse-Train, and their performance is reported on MMEVerse-Bench.
@@
TABLE XI: An example of multimodal emotion reasoning comparing our Emotion-LLaMAv2 with other MLLMs. Incorrect reasoning is marked in red, correct reasoning in blue, and hallucinations in green.
@@
Ablation on Modality Encoders. We conduct ablation studies to identify the optimal unimodal encoders and their multimodal combinations, with results shown in Table X. In the audio domain, HuBERT achieves the highest performance (58.92%), outperforming Whisper by 2.64%. For static image encoding, EVA proves superior (57.7%), which we attribute to its pre-alignment with the LLaMA2 weights used in our base model MiniGPTv2. For temporal video encoding, applying strong image encoders (e.g., EVA, CLIP) to individual frames is more effective than dedicated video encoders such as VideoMAE, as this approach better preserves fine-grained visual clues critical for emotion recognition.
@@ 5.6 Qualitative Analysis and Visualization
Qualitative Analysis of Emotion Reason. To illustrate the qualitative performance of Emotion-LLaMAv2, we present a comparison of emotion reasoning results across models. Table XI reports the reasoning outputs of several mainstream models. The video depicts a person smiling while questioning another individual, an expression of dissatisfaction suggesting an angry emotional state. Accurate reasoning for this sample requires integrating information from multiple modalities. PandaGPT and Valley capture the correct visual features but fail to incorporate other modalities, misclassifying the emotion as happy. Emotion-LLaMA reaches the correct inference, but its reasoning is compromised by hallucinations. AffectGPT provides a reasonable description yet misclassifies the emotion category. In contrast, Emotion-LLaMAv2 demonstrates a comprehensive understanding by recognizing the speaker’s tone and combining subtle facial expressions with multimodal information for accurate reasoning. This example highlights the superiority of our model in integrating emotional clues across modalities, yielding precise and contextually relevant recognition.

### v2

@@ 5.3 Main Results
TABLE VI: Overall performance comparison ($\%$) on recent MME-Emotion [86]. For both open-source and closed-source models, the best results are bold, and the second-best are highlighted in gray.
 Model LLM Size Avg Step Avg Token Rec-S Rea-S CoT-S
Open-source MLLMs
Qwen2-Audio [12] 7B 3.0 40.3 34.1 50.4 42.3
Audio-Reasoner [78] 7B 5.0 356.8 38.1 71.6 54.8
Qwen2.5-VL-7B [2] 7B 4.8 169.7 28.4 64.8 46.6
Qwen2.5-VL-72B [2] 72B 4.8 266.3 31.3 75.7 53.5
QVQ [68] 72B 5.5 899.9 31.4 70.1 50.8
Video-LLaVA [47] 7B 2.3 19.1 25.8 32.8 29.3
Video-LLaMA [88] 7B 4.5 122.5 26.1 48.5 37.3
Video-LLaMA2 [9] 7B 2.6 37.7 29.2 27.7 28.4
Qwen2.5-Omni [79] 7B 3.7 78.6 17.4 59.3 38.4
Emotion-LLaMA [7] 7B 1.0 2.3 25.1 0.4 12.8
HumanOmni [91] 7B 1.0 1.3 36.0 0.3 18.1
R1-Omni [90] 0.5B 5.0 156.2 26.3 58.6 42.4
AffectGPT [43] 7B 4.9 122.8 11.9 50.6 31.2
Emotion-LLaMAv2 (Stage 2) 7B 5.8 184.5 43.26 71.63 57.45
Closed-source MLLMs
GPT-4o [31] 4.4 169.4 27.8 79.8 53.8
GPT-4.1 [60] 5.2 141.2 28.8 65.2 47.0
Gemini-2.0-Flash [20] 4.1 64.7 36.3 60.0 48.1
Gemini-2.5-Flash [21] 4.3 261.8 34.7 52.7 43.7
Gemini-2.5-Pro [22] 5.1 538.6 39.3 72.7 56.0
TABLE VII: The Emotion-LLaMAv2 performance trained on different MER instruction tuning datasets.
@@
TABLE VIII: Ablation study of the Conv-Attention module.
@@
TABLE IX: Comparison of different pre-fusion architectures.
@@
TABLE X: Comparison of different training schemes.
@@
Results on MME-Emotion. MME-Emotion [86] is a very recent holistic evaluation benchmark designed for emotion recognition and reasoning. In Table VI, we present a comparison of our model, Emotion-LLaMAv2, with both open-source and closed-source multimodal LLMs using the standard evaluation metric from the benchmark. Remarkably, our model outperforms all others in both the recognition score (Rec-S) and Chain-of-Thought score (CoT-S). Additionally, it ranks second in the reasoning score (Rea-S) among open-source models. Notably, all closed-source multimodal LLMs rely solely on visual and textual information yet maintain competitive performance. Our state-of-the-art results on MME-Emotion indicate that advanced training on a large, multimodal emotion understanding dataset can surpass even commercialized large models trained on extensive datasets.
@@ 5.4 Data Evaluation
Effectiveness of MMEVerse. Table VII presents an evaluation of recent multimodal emotion instruction tuning datasets, including MERR-Coarse, MERR-Fine, and MER-Caption+. We train Emotion-LLaMAv2 on these datasets with their emotion categories and showcase its performance on both MER-UniBench and MMEVerse-Bench. The results reveal that training on MMEVerse-Train yields significantly better performance, particularly on MMEVerse-Bench. This can be attributed to two main factors: its data diversity and annotation reliability. On the one hand, its diverse sources, spanning from YouTube videos to interviews, capture more authentic "in-the-wild" emotional expressions, equipping the model with superior generalization capabilities. On the other hand, MMEVerse is more accurate in annotation quality since each emotion label is manually annotated, enabling the model to learn true multimodal-emotion correlations without label noise.
@@ 5.5 Ablation Studies
Effectiveness of the Conv-Attention Module. Table VIII reports the evaluation of the Conv-Attention module. It significantly improves Stage 1 performance, increasing the average score by 1.48% on MER-UniBench and 0.92% on MMEVerse-Bench. This highlights its importance: pre-fusion representations complement unimodal embeddings through interaction before feeding into the LLM. For models trained in Stage 2, the gains remain but are slightly reduced, as complex reasoning training can offset some improvements.
To further assess the pre-fusion module, we compare our Conv-Attn architecture against the Q-Former and the MLP-based attention mechanism from AffectGPT [43]. As shown in Table IX, the inclusion of any pre-fusion module consistently improves performance over the baseline, validating the strategy of refining modal features prior to joint processing by the LLM. Among individual architectures, both the Q-Former and our Conv model perform strongly, suggesting that preserving local temporal information during pre-fusion is particularly effective. Ultimately, Conv-Attn achieves the highest performance, reflecting a synergistic effect between its components: convolutional layers excel at extracting local temporal features, while attention mechanisms capture long-range dependencies and global context. By integrating these complementary strengths, the model produces richer and more comprehensive representations, leading to superior final performance.
Effectiveness of the perception-to-cognition training scheme. Table X presents an evaluation of different strategies for instruction tuning. While one might argue that a joint training scheme for recognition and reasoning is simpler and more efficient, akin to our preliminary Emotion-LLaMA, surprisingly, this straightforward approach yields significantly poorer performance on both benchmarks listed in Table X. This can be attributed to the difficulty of learning both perception and cognition abilities from scratch, in contrast to a curriculum learning approach.
TABLE XI: Ablation study on Modality Encoders. All the models are trained on MMEVerse-Train, and their performance is reported on MMEVerse-Bench.
@@
TABLE XII: An example of multimodal emotion reasoning comparing our Emotion-LLaMAv2 with other MLLMs. Incorrect reasoning is marked in red, correct reasoning in blue, and hallucinations in green.
@@
Ablation on Modality Encoders. We conduct ablation studies to identify the optimal unimodal encoders and their multimodal combinations, with results shown in Table XI. In the audio domain, HuBERT achieves the highest performance (58.92%), outperforming Whisper by 2.64%. For static image encoding, EVA proves superior (57.7%), which we attribute to its pre-alignment with the LLaMA2 weights used in our base model MiniGPTv2. For temporal video encoding, applying strong image encoders (e.g., EVA, CLIP) to individual frames is more effective than dedicated video encoders such as VideoMAE, as this approach better preserves fine-grained visual clues critical for emotion recognition.
@@ 5.6 Qualitative Analysis and Visualization
Qualitative Analysis of Emotion Reason. To illustrate the qualitative performance of Emotion-LLaMAv2, we present a comparison of emotion reasoning results across models. Table XII reports the reasoning outputs of several mainstream models. The video depicts a person smiling while questioning another individual, an expression of dissatisfaction suggesting an angry emotional state. Accurate reasoning for this sample requires integrating information from multiple modalities. PandaGPT and Valley capture the correct visual features but fail to incorporate other modalities, misclassifying the emotion as happy. Emotion-LLaMA reaches the correct inference, but its reasoning is compromised by hallucinations. AffectGPT provides a reasonable description yet misclassifies the emotion category. In contrast, Emotion-LLaMAv2 demonstrates a comprehensive understanding by recognizing the speaker’s tone and combining subtle facial expressions with multimodal information for accurate reasoning. This example highlights the superiority of our model in integrating emotional clues across modalities, yielding precise and contextually relevant recognition.